OpenStudy (anonymous):

what is the difference 9/4 - 1/7 -3/8 59/28 -8/3 67/28

OpenStudy (anonymous):

\[\frac{ 9 }{ 4 }-\frac{ 1 }{ 7 }=\frac{ 63 }{ 28 }-\frac{ 4 }{ 28 }=\frac{ 59 }{ 28 }=2\frac{ 3 }{ 28 }\]

OpenStudy (anonymous):

that isnt even a answer choice?

OpenStudy (anonymous):

2 3/28 = 59 /28

OpenStudy (anonymous):

there is \[\frac{ 59 }{ 28 }\]